OpenStudy (kaybabyluvzyhu16.):

I WILL FAN AND MEDAL YOU!!! Is anyone REALLY good at algebra? I got a super hard long equation i need help w

OpenStudy (danieldjpon3):

I am

OpenStudy (danieldjpon3):

I'm ready let'z go!

OpenStudy (kaybabyluvzyhu16.):

Okay thanks so much let me write it out :)

OpenStudy (danieldjpon3):

ok brb in ten mins, @danieldjpon3 when you done

OpenStudy (user123):

im redy letz go! h

OpenStudy (kaybabyluvzyhu16.):

\[(3^{8} * 2^{-5} * 9^{0})^{-2} *\left( 2^{-2} \right)/3{^{3}} )^{4} * 3^{28}\]

OpenStudy (danieldjpon3):

ok

OpenStudy (kaybabyluvzyhu16.):

(38 ⋅ 2−5 ⋅ 90)−2 ⋅ 2 to the power of negative 2 over 3 to the power of 3, whole to the power of 4 ⋅ 328 this is it n written form

OpenStudy (kaybabyluvzyhu16.):

i have to find the value

OpenStudy (danieldjpon3):

the answer or value of what?

OpenStudy (kaybabyluvzyhu16.):

the expression

OpenStudy (danieldjpon3):

(38*2 - 5*90) - 2 * ( 2^-2 / 3^3 ) - 4* 3^28 ( 190 - 450 ) - 2 * (0.25/27) - 9.15^13 -260 * -2 * (0.25/27 ) - 9.15 ^ 13 (-260 * -0.02) - 9.15 ^13 5.2 - 9.15^ 13 -3 . 15^12 I hope that helps!!!

OpenStudy (danieldjpon3):

well... @kaybabyluvzyhu16

OpenStudy (danieldjpon3):

???

jhonyy9 (jhonyy9):

@danieldjpon3 the first 3 is on exponent of 8 not 3 time 8 and these all time 2 on exponent of minus 5 time 9 on exponent of zero and all these inside parentheses on exponent of minus 2 - this is the first part of this exercise inside first parentheses so these all time 2 on exponent of minus 2 divide 3 on exponent of 3 @KayBabyLuvzYhu16.and after these there is one parentheses and exponent of 4 but i think that you missed the pair of this parenthese so bc. i dont see it nowhere please clarify and this time 3 on exponent of 28 are these above wrote correct in this way please ?

OpenStudy (campbell_st):

well the 1st thing you need to know is that \[a^0 = 1\] or any base to a power of zero is 1 so that should simplify the 1st set of brackets next, still looking at the 1st set of brackets apply the index law for power of a power \[(a^x)^y = a^{x \times y}\] so you multiply the power of -2 \[(3^8 \times 2^{-5})^{-2} = (3^8)^{-2} \times (2^{-5]})^{-2} \] in the next part of the equation re-write it as just simplify the power on the right hand side of the equation. apply the same rule to the next part of the \[\frac{2^{-2}}{3^4} = 2^{-2} \times 3^{-4}\] fractions can be written with negative powers s you now need to use the power of a power rule.... for \[(2^{-2} \times 3^{-4})^4 \] then the solution is simply a case of adding the powers of 3 and adding the powers of 2 to help get the final solution